OpenStudy (anonymous):

Find a unit vector in the direction of the given vector. w=-7i

zepdrix (zepdrix):

Take your vector w, and divide each component by it's length.\[\Large\rm \hat w=\frac{\vec w}{|w|}=\frac{<0,-7>}{7}=\left<\frac{0}{7},\frac{-7}{7}\right>=\left<0,-1\right>\]

zepdrix (zepdrix):

\(\Large\rm \hat w=-i\) Maybe I wrote that a little more complicated that I needed to :)

zepdrix (zepdrix):

The vector goes straight down, so it has a length of 7. So divide the vector by 7. Your unit vector is then -i

OpenStudy (anonymous):

ahh okay that's what I got, looks like my online homework is telling me the right answer is wrong again :-(. thanks

zepdrix (zepdrix):

Oh wait wait, is i supposed to be the x-component? like i,j ? or is it the imaginary unit? I guess I should have asked that first.

OpenStudy (anonymous):

yeah like i and j

zepdrix (zepdrix):

Oh then maybe they wanted the answer in component form? Like this? <-1,0>

zepdrix (zepdrix):

Sometimes those online things can be fussy >.<

OpenStudy (anonymous):

eh i doubt it, the other identical problems like it are in i and j :( oh well

zepdrix (zepdrix):

mm sucks :c